the length of a rectangle is 10 feet more than twice the width. the perimeter is 74. what is the length?
i know the answer is 28, but i don't know how to resolve it

Answer by ankor@dixie-net.com(22740)

the length of a rectangle is 10 feet more than twice the width.
the perimeter is 74. what is the length?
:
The perimeter equation
2L + 2W = 74
Simplify, divide by 2
L + W = 37
:
It says,"the length of a rectangle is 10 feet more than twice the width."
equation for this statement:
L = 2W + 10
:
Using the simplified perimeter equation, replace L with (2W+10)
2W + 10 + W = 37
3W = 37 - 10
3W = 27
W = 9
:
you can find the length: L = 2W + 10